Session: Oct 2011 (session 3)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

JIVE (Miguel Sanz at Thu Nov 10 05:34:42 2011) Minor failures:
Since scan nš 12,we rise the attenuation (in one dB more), becuse low elevation with bad weather conditions (cloudy an rainy), gave us too high temperatures.

Hartebeesthoek (Jonathan Quick at Thu Nov 10 05:41:46 2011) Success:
No known problems

Onsala (Michael Lindqvist at Thu Nov 10 07:23:03 2011) Success:
Light rain. Otherwise no known problem.

Yebes (Miguel Sanz at Thu Nov 10 08:17:59 2011) Success:
Sorry for using JIVE site instead Yebes.

Urumqi (Wenju Yang at Thu Nov 10 16:13:32 2011) Minor failures: HIGH
bad weather with little snow and thick fog. the LO of vc1 and vc2 was unlock at UT 23:30.

Effelsberg (Uwe Bach at Mon Nov 14 09:19:46 2011) Success:
Tsys for the BBC on Ori KL increased to 4000 K, but did not overflow.

Effelsberg (Uwe Bach at Mon Nov 14 09:23:22 2011) Success:
Sorry, Ori KL was EE008

Metsahovi (Rastorgueva at Mon Nov 14 14:55:55 2011) Success:
The session was carried out without any noticebla problems